# Towards Computation- and Communication-efficient Computational Pathology

### v1

@@ Methods
MAGA-GLTrans
MAGA-GLTrans is a computation-efficient, communication-efficient and storage-friendly WSI analysis pipeline. The same with existing mainstream WSI analysis pipelines, it consists of two main modules, a feature extractor (MAGA) and a predictor (GLTrans). The major difference is that MAGA uses low magnification inputs and the conventional encoders use high magnification ones.
Magnification Alignment (MAGA)
MAGA is a self-supervised training strategy that enables the expansion of an encoder, initially trained on high magnification images, to effectively handle low magnification inputs. The main concept of MAGA is to minimize the information gap between two magnification levels by aligning their respective feature spaces. To achieve this, two models with identical network structures are employed, including a well-trained feature extractor for high magnification level and a feature extractor for low magnification level (Figure 1b). $L1$ loss is applied to make the features extracted from different encoders as similar as possible, from the shallow layers to the deep layers.
$\mathcal{L}_{MAGA}=\sum_{i=1}^{n}{|X_{i}^{H}-X_{i}^{L}|}x$ (1)
where $X_{i}^{H}$ and $X_{i}^{L}$ denote the feature maps of the $i$-th block extracted from the high magnification encoder and the low magnification encoder, respectively. $n$ is the number of blocks. For example in ResNet50, $n$ is 5, where 1 to 4 correspond to layer1 to layer4 and 5 represents the output of the global average pooling layer. During MAGA training, we only update the low magnification encoder and freeze the high magnification one. The image resolution of the patches at different magnifications is 256$\times$256 (20×20$\times$20 ×), 128$\times$128 (10×10$\times$10 ×) and 64$\times$64 (5×5$\times$5 ×). The patches at the same location under different magnifications share exactly the same image content. For the WSIs without predefined specific magnification levels, we directly downsample the patches to the corresponding magnification level. Theoretically, using as large a dataset as possible to train a model can yield a diverse and robust feature extractor. Nonetheless, due to the constraints on computational resources, we train a distinct model for each dataset. 3,293,000, 12,244,980 and 2,821,910 non-overlapping histopathological image patches are extracted from TCGA-NSCLC, TCGA-Glioma and CAMELYON16 are used to train MAGA, respectively. The model parameters are updated using the Adam optimizer with an initial learning rate of 0.0001. The learning strategy is linear decay with a decay coefficient of 0.9. Since MAGA is trained in a self-supervised manner, the training procedure is performed without any manual intervention or labeling burden. In addition, the model-agnostic nature allows MAGA to be applied to any well-trained encoder, for example ImageNet pre-trained ResNet, histopathology-specific encoder CTransPath or other pathology foundation models.
@@
We used Python (v3.10.0) and Pytorch (v2.1.2) to build all deep learning models, experiments and analyses in this study. The ImageNet pre-trained weights of ResNet50 (https://download.pytorch.org/models/resnet50-0676ba61.pth) and histopathology images pre-trained weights of CTransPath [51] (https://github.com/Xiyue-Wang/TransPath) and UNI [23] (https://github.com/mahmoodlab/UNI) were used to train MAGA. The source codes of CLAM (https://github.com/mahmoodlab/CLAM) and TransMIL (https://github.com/szc19990412/TransMIL) were downloaded from the corresponding GitHub repositories. All the models, including MAGA-GLTrans, CLAM and TransMIL, as well as the downstream experiments were trained and conducted on a server equipped with a single NVIDIA GeForce RTX 4090 GPU with 256 GB of running memory. Reading of WSIs was done using the openslide-python (v1.3.1) and pyvips (v2.2.1), and writing of the images was done using Opencv-python (v4.8.1). We used the k-mean algorithm in Scikit-learn (v1.3.2) to perform patch-level clustering. The box plots, radar plots, heatmaps, histograms, scatter plots, ROC curves and the patch attention heatmap were drawn by the Python matplotlib (v3.8.2).
@@
MAGA-GLTrans can be accessed at https://github.com/Bingchao-Zhao/MAGA-GLTrans, including the model weights of MAGA and GLTrans for each specific task and their relevant source code.
@@
Extended Figure 1: Classification performance, computational efficiency and storage burden (10×10$\times$10 × versus 20×20$\times$20 ×, TCGA-NSCLC, TCGA-Glioma, CAMELYON16). a-f, the Classification performance of the 10-fold cross-validation experiment on three CPath tasks, including subtyping on TCGA-NSCLC (a, d), IDH mutation prediction on TCGA-Glioma (b, e) and lymph node metastasis detection on CAMELYON16 (c, f). a-c, Mean AUC of MAGA-GLTrans (10×10$\times$10 ×), GLTrans (10×10$\times$10 × and 20×20$\times$20 ×), CLAM (10×10$\times$10 × and 20×20$\times$20 ×) and TransMIL (10×10$\times$10 × and 20×20$\times$20 ×). d-f, Box plots demonstrate the results of each fold in the 10-fold cross-validation by resampling 500 times. We compare MAGA-GLTrans (10×10$\times$10 ×) with two competitors at different magnification levels (10×10$\times$10 ×, top) and (20×20$\times$20 ×, bottom). g-l, Mean computational time (g, i, k) of each model and mean size of the WSI files (h, j, l) on three different datasets.
@@
Extended Figure 2: Classification performance, computational efficiency and storage burden (5×5$\times$5 × versus 20×20$\times$20 ×, TCGA-Glioma). a-d, Classification performance of the 10-fold cross-validation experiment on TERT mutation prediction (a, c) and 1p19q codeletion prediction (b, e) on TCGA-Glioma. a-b, Mean AUC of MAGA-GLTrans (5$\times$), GLTrans (5$\times$ and 20$\times$), CLAM (5$\times$ and 20$\times$) and TransMIL (5$\times$ and 20$\times$). c-d, Box plots demonstrate the results of each fold in the 10-fold cross-validation by resampling 500 times. We compare MAGA-GLTrans (5$\times$) with two competitors at different magnification levels (5$\times$, top) and (20$\times$, bottom). e-h, Mean computational time (e, g) of each model and mean size of the WSI files (f, h) in two different tasks.
@@
Extended Figure 3: Classification performance, computational efficiency and storage burden (10×10$\times$10 × versus 20×20$\times$20 ×, TCGA-Glioma). a-d, Classification performance of the 10-fold cross-validation experiment on TERT mutation prediction (a, c) and 1p19q codeletion prediction (b, e) on TCGA-Glioma. a-b, Mean AUC of MAGA-GLTrans (10×10$\times$10 ×), GLTrans (10×10$\times$10 × and 20×20$\times$20 ×), CLAM (10×10$\times$10 × and 20×20$\times$20 ×) and TransMIL (10×10$\times$10 × and 20×20$\times$20 ×). c-d, Box plots demonstrate the results of each fold in the 10-fold cross-validation by resampling 500 times. We compare MAGA-GLTrans (10×10$\times$10 ×) with two competitors at different magnification levels (10×10$\times$10 ×, top) and (20×20$\times$20 ×, bottom). e-h, Mean computational time (e, g) of each model and mean size of the WSI files (f, h) in two different tasks.
@@
Extended Figure 4: Visualization of the features extracted by MAGA (CAMELYON16). a, The prediction results (top) of the test set on the breast cancer lymph node metastasis detection task (CAMELYON16, $n=129$) under different magnification levels. Patients are sorted by the ground truth labels. Heatmaps (bottom) demonstrate the similarity of the features between high magnification level and low magnification level (with and without MAGA). We show the feature similarity in slide level, which is the mean of all the patches. Red represents higher similarity. b, From top left to bottom right: a sample case WSI of metastasis, visualization of the unsupervised clustering of the patch-level features (top right) and the heatmap (bottom right) at 20×20$\times$20 ×, 5×5$\times$5 × with MAGA and 5×5$\times$5 × without MAGA.

### v2

@@ Methods
MAG-GLTrans
MAG-GLTrans is a computation-efficient, communication-efficient and storage-friendly WSI analysis pipeline. The same with existing mainstream WSI analysis pipelines, it consists of two main modules, a feature extractor (MAG) and a predictor (GLTrans). The major difference is that MAG uses low magnification inputs and the conventional encoders use high magnification ones.
Magnification Alignment (MAG)
MAG is a self-supervised training strategy that enables the expansion of an encoder, initially trained on high magnification images, to effectively handle low magnification inputs. The main concept of MAG is to minimize the information gap between two magnification levels by aligning their respective feature spaces. To achieve this, two models with identical network structures are employed, including a well-trained feature extractor for high magnification level and a feature extractor for low magnification level (Figure 1b). $L1$ loss is applied to make the features extracted from different encoders as similar as possible, from the shallow layers to the deep layers.
$\mathcal{L}_{MAG}=\sum_{i=1}^{n}{|X_{i}^{H}-X_{i}^{L}|}x$ (1)
where $X_{i}^{H}$ and $X_{i}^{L}$ denote the feature maps of the $i$-th block extracted from the high magnification encoder and the low magnification encoder, respectively. $n$ is the number of blocks. For example in ResNet50, $n$ is 5, where 1 to 4 correspond to layer1 to layer4 and 5 represents the output of the global average pooling layer. During MAG training, we only update the low magnification encoder and freeze the high magnification one. The image resolution of the patches at different magnifications is 256$\times$256 (20×20$\times$20 ×), 128$\times$128 (10×10$\times$10 ×) and 64$\times$64 (5×5$\times$5 ×). The patches at the same location under different magnifications share exactly the same image content. For the WSIs without predefined specific magnification levels, we directly downsample the patches to the corresponding magnification level. Theoretically, using as large a dataset as possible to train a model can yield a diverse and robust feature extractor. Nonetheless, due to the constraints on computational resources, we train a distinct model for each dataset. 3,293,000, 12,244,980 and 2,821,910 non-overlapping histopathological image patches are extracted from TCGA-NSCLC, TCGA-Glioma and CAMELYON16 are used to train MAG, respectively. The model parameters are updated using the Adam optimizer with an initial learning rate of 0.0001. The learning strategy is linear decay with a decay coefficient of 0.9. Since MAG is trained in a self-supervised manner, the training procedure is performed without any manual intervention or labeling burden. In addition, the model-agnostic nature allows MAG to be applied to any well-trained encoder, for example ImageNet pre-trained ResNet, histopathology-specific encoder CTransPath or other pathology foundation models.
@@
We used Python (v3.10.0) and Pytorch (v2.1.2) to build all deep learning models, experiments and analyses in this study. The ImageNet pre-trained weights of ResNet50 (https://download.pytorch.org/models/resnet50-0676ba61.pth) and histopathology images pre-trained weights of CTransPath [51] (https://github.com/Xiyue-Wang/TransPath) and UNI [23] (https://github.com/mahmoodlab/UNI) were used to train MAG. The source codes of CLAM (https://github.com/mahmoodlab/CLAM) and TransMIL (https://github.com/szc19990412/TransMIL) were downloaded from the corresponding GitHub repositories. All the models, including MAG-GLTrans, CLAM and TransMIL, as well as the downstream experiments were trained and conducted on a server equipped with a single NVIDIA GeForce RTX 4090 GPU with 256 GB of running memory. Reading of WSIs was done using the openslide-python (v1.3.1) and pyvips (v2.2.1), and writing of the images was done using Opencv-python (v4.8.1). We used the k-mean algorithm in Scikit-learn (v1.3.2) to perform patch-level clustering. The box plots, radar plots, heatmaps, histograms, scatter plots, ROC curves and the patch attention heatmap were drawn by the Python matplotlib (v3.8.2).
@@
MAG-GLTrans can be accessed at https://github.com/Bingchao-Zhao/MAG-GLTrans, including the model weights and the relevant source code.
@@
Extended Figure 1: Classification performance, computational efficiency and storage burden (10×10$\times$10 × versus 20×20$\times$20 ×, TCGA-NSCLC, TCGA-Glioma, CAMELYON16). a-f, the Classification performance of the 10-fold cross-validation experiment on three CPath tasks, including subtyping on TCGA-NSCLC (a, d), IDH mutation prediction on TCGA-Glioma (b, e) and lymph node metastasis detection on CAMELYON16 (c, f). a-c, Mean AUC of MAG-GLTrans (10×10$\times$10 ×), GLTrans (10×10$\times$10 × and 20×20$\times$20 ×), CLAM (10×10$\times$10 × and 20×20$\times$20 ×) and TransMIL (10×10$\times$10 × and 20×20$\times$20 ×). d-f, Box plots demonstrate the results of each fold in the 10-fold cross-validation by resampling 500 times. We compare MAG-GLTrans (10×10$\times$10 ×) with two competitors at different magnification levels (10×10$\times$10 ×, top) and (20×20$\times$20 ×, bottom). g-l, Mean computational time (g, i, k) of each model and mean size of the WSI files (h, j, l) on three different datasets.
@@
Extended Figure 2: Classification performance, computational efficiency and storage burden (5×5$\times$5 × versus 20×20$\times$20 ×, TCGA-Glioma). a-d, Classification performance of the 10-fold cross-validation experiment on TERT mutation prediction (a, c) and 1p19q codeletion prediction (b, e) on TCGA-Glioma. a-b, Mean AUC of MAG-GLTrans (5$\times$), GLTrans (5$\times$ and 20$\times$), CLAM (5$\times$ and 20$\times$) and TransMIL (5$\times$ and 20$\times$). c-d, Box plots demonstrate the results of each fold in the 10-fold cross-validation by resampling 500 times. We compare MAG-GLTrans (5$\times$) with two competitors at different magnification levels (5$\times$, top) and (20$\times$, bottom). e-h, Mean computational time (e, g) of each model and mean size of the WSI files (f, h) in two different tasks.
@@
Extended Figure 3: Classification performance, computational efficiency and storage burden (10×10$\times$10 × versus 20×20$\times$20 ×, TCGA-Glioma). a-d, Classification performance of the 10-fold cross-validation experiment on TERT mutation prediction (a, c) and 1p19q codeletion prediction (b, e) on TCGA-Glioma. a-b, Mean AUC of MAG-GLTrans (10×10$\times$10 ×), GLTrans (10×10$\times$10 × and 20×20$\times$20 ×), CLAM (10×10$\times$10 × and 20×20$\times$20 ×) and TransMIL (10×10$\times$10 × and 20×20$\times$20 ×). c-d, Box plots demonstrate the results of each fold in the 10-fold cross-validation by resampling 500 times. We compare MAG-GLTrans (10×10$\times$10 ×) with two competitors at different magnification levels (10×10$\times$10 ×, top) and (20×20$\times$20 ×, bottom). e-h, Mean computational time (e, g) of each model and mean size of the WSI files (f, h) in two different tasks.
@@
Extended Figure 4: Visualization of the features extracted by MAG (CAMELYON16). a, The prediction results (top) of the test set on the breast cancer lymph node metastasis detection task (CAMELYON16, $n=129$) under different magnification levels. Patients are sorted by the ground truth labels. Heatmaps (bottom) demonstrate the similarity of the features between high magnification level and low magnification level (with and without MAG). We show the feature similarity in slide level, which is the mean of all the patches. Red represents higher similarity. b, From top left to bottom right: a sample case WSI of metastasis, visualization of the unsupervised clustering of the patch-level features (top right) and the heatmap (bottom right) at 20×20$\times$20 ×, 5×5$\times$5 × with MAG and 5×5$\times$5 × without MAG.